OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = -3x + 9

OpenStudy (anonymous):

Hi need help correcting y answers

OpenStudy (anonymous):

Slope -3, y intercept 9.

OpenStudy (anonymous):

Cool i got the same one

OpenStudy (anonymous):

Can you help mew with 4 more?

OpenStudy (anonymous):

Good 4 you buddie :]

OpenStudy (anonymous):

Just help me correct my answer

OpenStudy (anonymous):

Sure thang

OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = x - 3

OpenStudy (anonymous):

is it Slope is 0 and Y intercept is -3

OpenStudy (anonymous):

No is it the slop is 1 and the intercept is -3?

OpenStudy (anonymous):

Yessir

OpenStudy (anonymous):

slope: 3, y-intercept: 1 slope: 3, y-intercept: 1 slope: 1, y-intercept: 3 slope: 1, y-intercept: 3

OpenStudy (anonymous):

The buttom one is a -1 and the intercept just a regular -3

OpenStudy (anonymous):

So is it Slope is 1 and Y intercept is -3

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Ok 3 moire

OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = 4x + 4 (Points : 1) slope: 4, y-intercept: 0 slope: 4, y-intercept: 4 slope: 0, y-intercept: 0 slope: 0, y-intercept: 4

OpenStudy (anonymous):

Num 2

OpenStudy (anonymous):

Is it the Slope is 4 and the Y intercept is 4?

OpenStudy (anonymous):

Ye

OpenStudy (anonymous):

So its B right :D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok 2 more!

OpenStudy (anonymous):

Transform this equation to slope-intercept form: 6x + 4y = 8. (Points : 1)

OpenStudy (anonymous):

I akm sovling it XD

OpenStudy (anonymous):

y=-1/2 + 2

OpenStudy (anonymous):

This is my answer is it correct?

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 3 }x+2\]

OpenStudy (anonymous):

The answer isnt one of those

OpenStudy (anonymous):

Let me screencp it!

OpenStudy (anonymous):

The one we are on is 3

OpenStudy (anonymous):

It's c

OpenStudy (anonymous):

Its c?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

C as y=-3/2+2

OpenStudy (anonymous):

Yes, you move x over, then divide y.

OpenStudy (anonymous):

Real quick what is -1.5 written as a fraction

OpenStudy (anonymous):

\[-1\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

how we get -3/2

OpenStudy (anonymous):